OpenStudy (anonymous):

Multiply and simplify:

OpenStudy (anonymous):

The problem is: \[2\sqrt{5}(3\sqrt{2}+3\sqrt{5})\]

OpenStudy (lilg132):

\[30 + 6\sqrt{10}\]

jimthompson5910 (jim_thompson5910):

\[2\sqrt{5}(3\sqrt{2}+3\sqrt{5})\] \[2\sqrt{5}(3\sqrt{2})+2\sqrt{5}(3\sqrt{5})\] \[2*3\sqrt{5*2}+2*3\sqrt{5*5}\] \[6\sqrt{10}+6\sqrt{25}\] \[6\sqrt{10}+6*5\] \[6\sqrt{10}+30\] \[30+6\sqrt{10}\]

OpenStudy (anonymous):

The answer choices: A. \[15\sqrt{10}+6\sqrt{5}\] B. \[6\sqrt{10}+6\sqrt{5}\] C. \[6\sqrt{10}+30\]

OpenStudy (lilg132):

c

jimthompson5910 (jim_thompson5910):

yes choice C, you can rearrange the terms in any order